OpenStudy (anonymous):

Can someone tell me how to solve this? -3v=195 A. -45 B. 45 C. 65 D.-65 @whpalmer4 @zepdrix @MathNerd67

OpenStudy (anonymous):

divide by -3

OpenStudy (jdoe0001):

"divide by -3" <--- both sides that is

OpenStudy (whpalmer4):

\[-3v=195\]You want \[v = \]The only thing standing between you and what you want is that pesky \(-3\) in front of \(v\). Both sides of \(-3v = 195\) are equal, so they will both still be equal if you divide both sides by \(-3\), and that will give you \(v=\) just as you want. \[-3v = 195\]\[\frac{-3v}{-3} = \frac{195}{-3}\]\[\frac{\cancel{-3}v}{\cancel{-3}} = \frac{195}{-3}\]\[v = \frac{195}{-3} = \]